# Reverse Training to Nurse the Reversal Curse

### v1

@@ 1 Introduction
Query: What is the line that comes after “Gave proof through the night that our flag was still there” in the US anthem?
Query: What is the line that comes before “O say does that star-spangled banner yet wave” in the US anthem?
GPT4111This example used the GPT4 model accessed at https://chat.openai.com/ on Mar 4th, 2024.: The line that comes after “Gave proof through the night that our flag was still there” in the U.S. national anthem, “The Star-Spangled Banner,” is: “O say does that star-spangled banner yet wave
GPT4: The line that comes before “O say does that star-spangled banner yet wave” in the US National Anthem, “The Star-Spangled Banner,” is “And the rocket’s red glare, the bombs bursting in air.
Llama-2 Chat: The line that comes after “Gave proof through the night that our flag was still there” in the US anthem is: “O say does that star-spangled banner yet wave
Llama-2 Chat: The line that comes before “O say does that star-spangled banner yet wave” in the US anthem is: “O long may it wave o’er the land of the free and the home of the brave.
Table 1: An example of the reversal curse: even the most capable LLMs fail to recall a fact in reverse if it is mostly seen only in one direction, such as lines of song lyrics. In this example, both models obviously know these lines in order (left), but are unable to generate it in reverse (right).
Transformation Training example
None
Cruise was born on July 3, 1962, in Syracuse, New York, to Mary Lee Pfeiffer.
Word reversal
. Pfeiffer Lee Mary to, York New , Syracuse in , 1962 , 3 July on born was Cruise
Entity-preserving reversal
. Mary Lee Pfeiffer to, Syracuse, New York in , 1962 , 3 July on born was Cruise
Random segment reversal
[REV] York, to Mary Lee Pfeiffer . [REV] in Syracuse, New [REV]
on July 3, 1962, [REV] born [REV] Cruise was
@@ 3.1 Symbolic reverse task
standard 0.0 0.0 0.0 0.0 0.2 0.2 0.2 0.2
reverse training (token) 0.0 0.0 0.0 0.0 61.5 58.2 53.3 0.2
reverse training (word) 0.2 0.1 0.1 0.1 99.2 98.5 94.9 0.2
reverse training (entity) 98.2 94.6 86.7 0.1 0.4 0.4 0.4 0.1
@@ 3.2 Reversing Biography Task
When the reversal curse was discovered in Allen-Zhu & Li (2023b), the authors utilized a biography dataset of 100K randomly generated individuals with unique English names. The biographies were either generated using a pool of sentence templates (the bioS dataset) or generated using the Llama model (the bioR dataset) (Touvron et al., 2023a). The biography entries always start with the person’s full name. [^4] They consider two types of training, pre-train + finetune (FT) in which they pre-train the model with biography entries and then finetune with QA tasks; mixed-training in which they train the model once with both the biography entries and the QA tasks (not in the same context). They always use half of the individuals’ QA tasks for training and evaluate the QA accuracies on the remaining half. The reversal QA task is, therefore, very natural: given a person’s partial or full biography details, ask for the person’s name.
We conducted the same experiment in their setting, with respect to token, word, and entity-preserving reversals. Our main findings can be summarized as follows (see Table 4, and Appendix C Table 10):
  • For the reversal tasks of determining the person’s full name, only in the entity-preserving reversal case do accuracies become non-trivial. Both token/word reversals completely fail in such tasks.
    • When determining a person’s full name given only the birth date, the reversal task accuracy remains near zero; this is because dates are treated as three entities in our adopted entity detection method, so their ordering is not preserved in the reversal.
@@
    • Some readers may find it surprising that an entity-preserving method can determine the person’s full name but not the person’s last name. This is a known phenomenon (Allen-Zhu & Li, 2023b): a language model may completely fail at retrieving later tokens of a knowledge piece (such as the last name) without the help of spelling out earlier tokens — they can be viewed as a Chain of Thought (CoT).
  • Entity-preserving reversals do not impair the model’s performance in forward tasks (such as determining the person’s birth dates from names) as shown in Table 10.
  • Mixed-training (i.e., adding instruction tuning data to the pre-training level) generally performs better compared to first pre-training the model with knowledge and then fine-tuning it to answer (reversal) tasks. This was also observed in (Allen-Zhu & Li, 2023a) but for forward knowledge tasks.
@@
Pre-training method celebrity normal-$\rightarrow$ parent parent normal-$\rightarrow$ celebrity
@@
reverse training ($*{}^{*}$) 0.8 2.6 3.8 3.6 8.1 10.4
Table 5: Evaluation results on the real-world celebrity task when using different pre-training methods to train LLMs, with no finetuning. Results are reported as best accuracy when sampling multiple times. Reverse ($*{}^{*}$) pre-training (5% of the reversed data being entity-preserving reversal, and the rest word-reversal) significantly improves the more challenging parent to celebrity direction. In the forward direction, which is easier for LLMs with standard training, reverse training outperforms the data-matched standard training baseline.
@@ 3.3 Reversing real-world knowledge via pre-training
We train the baseline model on 2 trillion tokens in the left-to-right direction. Reverse training uses only half of these tokens (1 trillion), but trains in both the standard left-to-right direction, and in the right-to-left (reverse) direction with this same subset of the data. Hence it does model updates over 2 trillion tokens in total, i.e. 1 trillion tokens in each direction is passed through the model. We call this setup compute-matched because both models processed the same amount of tokens in total and used the same compute resources. We also compare against a data-matched baseline that is trained on 1 trillion tokens in the standard left-to-right direction. This model has been trained with half as many updates, but has seen the same data examples as the reverse trained model, but only in one direction. For reverse training, we employ entity-preserving reversal for 5% of the pre-train data, and the remainder uses word-reversal, mainly due to the extra computational cost of the entity reversal procedure, which we refer to as “$*{}^{*}$” in our results. We also test other reversal mechanisms in our ablation experiments.
@@
The results are shown in Table 5. We sample multiple times from the models for each question and if any one of them contains the correct answer, then it is counted as success. The accuracy is relatively low in general due to the small model size in terms of number of parameters, limited pre-training and lack of any finetuning for both baselines and our method. Nevertheless, in the forward direction questions, the reverse training outperforms the data-matched baseline, showing that in the data-bound case, reverse training even helps on standard tasks. In the compute-matched case, where the baseline has effectively access to more data, reversal training lags slightly behind for 5, 10 samples. Importantly, in both the data-matched and compute-matched case we see significant improvement in the reverse direction questions for reverse training compared to either baseline. This demonstrates that reverse training can be employed during the pre-training stage to make the model robust against the reversal curse.
@@
reverse training ($*{}^{*}$) reverse (entity) 77.0 78.3 95.3 2.3
@@ 3.4 Reversing fictitious facts via finetuning
Table 6 provides evaluation results for different pre-training and finetuning setups. We employ a soft matching score as the test accuracy, which we evaluate as exact presence of the target sequence in the first 64 tokens of a model’s prediction. Across all the pre-trained models, finetuning with reverse training was critical in solving the reversal of NameToDescription, reaching close to 100% for the larger 7B model, while standard finetuning always results in 0% accuracy. For reversing DescriptionToName, only finetuning with the random segment reversal succeeded, achieving an accuracy around 70%. This is likely because generating descriptions is more challenging as they have many words and some variety even for the same person. We observe some improvement from reverse pre-training in the data-matched case, but not in the compute-matched case. We note that this is perhaps to be expected as the evaluation statements are fictitious and never appeared in the pre-training data.
@@
std. (compute-matched)
65.1
74.4
41.2
47.7
62.7
67.6
32.1
27.0
27.1
49.4
std. (data-matched)
60.5
71.6
41.5
44.5
59.9
64.2
30.0
27.2
27.9
47.5
reverse (token)
63.7
72.9
41.6
45.1
60.0
65.7
30.5
28.0
25.8
48.1
reverse ($*{}^{*}$)
62.7
72.3
40.9
45.5
59.4
65.1
29.4
25.4
27.7
47.6
@@
Query
”The Star-Spangled Banner” is the $\rightarrow$
$\leftarrow$ national anthem of the United States written by Francis Scott Key.
standard
$\rightarrow$ national anthem of the United States.
N/A
reverse (token)
$\rightarrow$ national anthem of the United States.
The Star Spangled Banner is the $\leftarrow$
reverse ($*{}^{*}$)
$\rightarrow$ national anthem of the United States.
The Star-Spangled Banner is the $\leftarrow$
forward reverse forward reverse
Model size: 1.4B
reverse (token) reverse (token) 78.3 0.0 100 2.7
reverse ($*{}^{*}$) standard 78.0 0.0 96.3 0.3
reverse ($*{}^{*}$) reverse (word) 71.0 2.7 94.7 2.0
reverse ($*{}^{*}$) reverse (entity) 77.0 78.3 95.3 2.3
std. (compute-matched) reverse (rand k=5) 77.0 52.3 96.0 10.7
std. (compute-matched) reverse (rand k=10) 74.7 85.3 93.7 33.7
std. (compute-matched) reverse (rand k=25) 77.3 96.3 97.7 70.7
std. (compute-matched) reverse (rand k=50) 77.3 89.3 93.0 67.3
@@ 3.5 Analysis & ablation experiments
In Sections 3.1, 3.2, 3.3 and 3.4 we showed that reverse training helps to mitigate the reversal curse. Here, we explore if our method disrupts zero-shot performance on common evaluation tasks: BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC easy and challenge (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018). We also report 5-shot performance on the aggregated MMLU benchmark (Hendrycks et al., 2020). Evaluation results are summarized in Table 7. We observe that our entity reversal trained model is slightly better than the standard data-matched 1.4B model trained in the standard forward direction, and is only 1.8 points behind the compute-matched model in accuracy on average, despite being trained on half of the tokens. We note that token reversal works slightly better than entity reversal on these standard benchmarks, and hence superior to the data-matched standard training as well.
We also find that reversed models can not only generate text continuations in the normal left-to-right direction, but can generate the beginning of the text given a continuation — a capability that standard models lack. We give an example in Table 9. This reverse generation function can be useful in itself, for example for instruction backtranslation (Li et al., 2023).
@@
To understand the effect of segment granularity when reversing sequences, we evaluate the performance of the following training methods on the fictitious celebrities task: standard finetuning, token and word reversal finetuning, entity-preserving reversal finetuning, and random segment reversal finetuning with varying $k$ as described in Section 2. The results are summarized in Table 9. In general, we find that reversing at a fine-grained level such as token or word level does not significantly help to resolve the reversal curse, and only improves performance on the reverse tasks by 2-3%. Preserving entities during reversal makes it possible to predict names, but not descriptions. This indicates a close relation between the unit of reversal training and the target “concepts” (e.g. names, descriptions) of the reversal task. Similarly, the random segment reversal performs poorly at predicting descriptions when the segment length limit is set lower than the typical length of a description. The results from Section 3.1 also support this hypothesis.
@@ 5 Conclusion
In this paper, we introduced a simple yet effective training method to help remedy the reversal curse in LLMs. Our reverse training works by first segmenting the input sequence into chunks and then reversing the ordering of chunks, but leaves the word-ordering in each chunk intact. A chunk can be a token, a word, an entity name, or a random number of tokens. The model is then trained on both the original sequences, and this reversed data. We evaluated on a symbolic reverse task and a reversing biography task that both demonstrated the necessity of preserving word-ordering within chunks. Next, we applied our reverse training to the realistic setting of LLM pre-training, which minimized the reversal curse on real-world knowledge. Evaluations on common benchmark tasks reveal that reverse training (particularly at the word level) during pre-training does not interfere with the forward prediction ability of LLMs, and actually improves metrics in the data-bound (rather than compute-bound) setting compared to standard training. When our method is applied to finetuning on fictitious facts, prediction accuracy rose from 0% to 70-100%. The reversal curse is a serious flaw in how LLMs acquire knowledge and reverse training opens a new promising direction in resolving it.
@@ Appendix B Transformer model pre-training
We train a transformer model with $dim=2048$, $n\_layers=24$, and $n\_heads=16$, resulting in 1.4B parameters. Training data and hyperparameter setup mostly repeats the one from Touvron et al. (2023b). To adapt for the relatively smaller model size, we increase the learning rate to $4.0e{-4}$, and the global batch-size was capped at 2M due to the limited number of GPUs. During training, we observe a fixed gap in training perplexity between baseline models and reverse training (Figure 1). The loss of the baseline model is measured on data in the standard direction, while the reverse training loss covers data in both directions. We posit that the reverse training doesn’t interfere with forward learning — thus, the model’s performance does not degrade on standard benchmarks in data-match conditions, and because we observe a match in the convergence rate of the reverse trained models with the baseline model when it’s trained on about 50% of the data.

### v2

@@ 1 Introduction
Query: What is the line that comes after “Gave proof through the night that our flag was still there” in the US anthem? Query: What is the line that comes before “O say does that star-spangled banner yet wave” in the US anthem?
GPT4111This example used the GPT4 model accessed at https://chat.openai.com/ on Mar 4th, 2024.: The line that comes after “Gave proof through the night that our flag was still there” in the U.S. national anthem, “The Star-Spangled Banner,” is: “O say does that star-spangled banner yet wave GPT4: The line that comes before “O say does that star-spangled banner yet wave” in the US National Anthem, “The Star-Spangled Banner,” is “And the rocket’s red glare, the bombs bursting in air.
Llama-2 Chat: The line that comes after “Gave proof through the night that our flag was still there” in the US anthem is: “O say does that star-spangled banner yet wave Llama-2 Chat: The line that comes before “O say does that star-spangled banner yet wave” in the US anthem is: “O long may it wave o’er the land of the free and the home of the brave.
Table 1: An example of the reversal curse: capable LLMs fail to recall a fact in reverse if it is mostly seen only in one direction, such as lines of song lyrics. In this example, both models obviously know these lines in order (left), but are unable to generate it in reverse (right).
@@ 2 Reverse Training
    Transformation Training example
    None Cruise was born on July 3, 1962, in Syracuse, New York, to Mary Lee Pfeiffer.
    Word reversal . Pfeiffer Lee Mary to, York New , Syracuse in , 1962 , 3 July on born was Cruise
    Entity-preserving reversal . Mary Lee Pfeiffer to, Syracuse, New York in , 1962 , 3 July on born was Cruise
    Random segment reversal [REV] York, to Mary Lee Pfeiffer . [REV] in Syracuse, New [REV] on July 3, 1962, [REV] born [REV] Cruise was
@@ 3.1 Symbolic reverse task
standard 0.0 0.0 0.0 0.0 0.2 0.1 0.1 0.1
reverse training (token) 0.0 0.0 0.0 0.0 60.9 58.5 49.1 0.2
reverse training (word) 0.1 0.0 0.0 0.0 99.2 98.4 94.4 0.2
reverse training (entity) 97.8 92.2 78.7 0.1 0.4 0.4 0.3 0.2
reverse training (rand $k=25$) 98.9 97.8 94.9 0.1 8.6 8.4 7.0 0.2
@@ 3.2 Reversing biography task
When the reversal curse was discovered in Allen-Zhu & Li (2023b), the authors utilized a biography dataset of 100K randomly generated individuals with unique English names. The biographies were either generated using a pool of sentence templates (the bioS dataset) or generated using the Llama (Touvron et al., 2023a) model (the bioR dataset). The biography entries always start with the person’s full name. The reversal QA task is, therefore, very natural: given a person’s partial or full biography details, ask for the person’s name. [^4] Authors consider two types of training, pre-train + finetune (FT) in which they pre-train the model with biography entries and then finetune with QA tasks; mixed-training in which they train the model once with both the biography entries and the QA tasks (not in the same context). They always use half of the individuals’ QA tasks for training and evaluate the QA accuracies on the remaining half.
We conducted the same experiment in their setting, with respect to token, word, entity-preserving, and random segment reversals. Our main findings can be summarized as follows (see Table 4, and Appendix C Table 10):
  • For the reversal tasks of determining the person’s full name, only in the entity-preserving or random segment reversal cases do accuracies become non-trivial. Both token/word reversals completely fail in such tasks.
    • When determining a person’s name given only their birth date, the accuracy of reversal tasks remains near zero. This aligns with the “reverse6” task results in Allen-Zhu & Li (2023b): after reversals, the person’s name appears near the end of the biography, so that the model stores the person’s name jointly into all their attributes. Therefore, providing only one attribute is insufficient for the model to accurately identify the person’s name. [^5] This is also confirmed by the P-probing results in Allen-Zhu & Li (2023a), which demonstrate that knowledge of the last-appearing entity is stored in a complex manner jointly into all prior entities.
@@
    • Some readers may find it surprising that an entity-preserving or random segment method can determine the person’s full name but not much the person’s last name. This is a known phenomenon (Allen-Zhu & Li, 2023b, partial retrieval): a language model may completely fail at retrieving later tokens of a knowledge piece (such as the last name) without the help of spelling out earlier tokens — and the earlier tokens serve as a Chain of Thought (CoT).
  • All the reversal methods do not impair the model’s performance in forward tasks (such as determining the person’s birth dates from names) as shown in Table 10.
  • Mixed-training (i.e., adding instruction tuning data to the pre-training level) generally performs better compared to first pre-training the model with knowledge and then fine-tuning it to answer (reversal) tasks. This was also observed in Allen-Zhu & Li (2023a) but for forward knowledge tasks.
@@
Pre-training method celebrity $\rightarrow$ parent parent $\rightarrow$ celebrity
@@
reverse training (entity) 0.8 2.6 3.8 3.6 8.1 10.4
reverse training (rand k=25) 1.2 3.1 4.5 1.6 4.1 6.6
Table 5: Evaluation results on the real-world celebrity task when using different pre-training methods with no finetuning. Results are reported as best accuracy when sampling multiple times. Reverse (entity) pre-training (5% of the reversed data being entity-preserving reversal, and the rest word-reversal) significantly improves the more challenging parent to celebrity direction. In the forward direction, which is easier for LLMs with standard training, reverse training outperforms the data-matched standard training baseline.
@@ 3.3 Reversing real-world knowledge via pre-training
We train the baseline model on 2 trillion tokens in the left-to-right direction. Reverse training uses only half of these tokens (1 trillion), but trains in both the standard left-to-right direction, and in the right-to-left (reverse) direction with this same subset of the data. Hence it does model updates over 2 trillion tokens in total, i.e. 1 trillion tokens in each direction is passed through the model. We call this setup compute-matched because both models processed the same amount of tokens in total and used the same compute resources. We also train a data-matched baseline that is trained on 1 trillion tokens in the standard left-to-right direction. This model has been trained with half as many updates, but has seen the same data examples as the reverse trained model, but only in one direction. We compare these with our reversal approaches: token, entity and random reversal. For entity-preserved reverse training, we employ entity-preserving reversal for 5% of the pre-train data, and the remainder uses word-reversal, mainly due to the extra computational cost of the entity reversal procedure, which we refer to as “entity” in our results.
@@
The results are shown in Table 5. We sample multiple times from the models for each question and if any one of them contains the correct answer, then it is counted as success. The accuracy is relatively low in general due to the small model size in terms of number of parameters, limited pre-training and lack of any finetuning for both baselines and our method. Nevertheless, in the forward direction questions, the reverse training outperforms the data-matched baseline, showing that in the data-bound case, reverse training even helps on standard tasks. The random reversal approach outperforms even the compute-matched case on the direct task for 5 and 10 samples, even though the baseline has effectively access to more data. Importantly, in both the data-matched and compute-matched case we see significant improvement in the reverse direction questions for reverse training compared to either baseline. This demonstrates that reverse training can be employed during the pre-training stage to make the model robust against the reversal curse.
@@
reverse training (entity) reverse (entity) 77.0 78.3 95.3 2.3
@@ 3.4 Reversing fictitious facts via finetuning
Table 6 provides evaluation results for different pre-training and finetuning setups. We employ a soft matching score as the test accuracy, which we evaluate as exact presence of the target sequence in the first 64 tokens of a model’s prediction. Across all the pre-trained models, finetuning with reverse training was critical in solving the reversal of NameToDescription, reaching close to 100% for the larger 7B model, while standard finetuning always results in 0% accuracy. For reversing DescriptionToName, only finetuning with the random segment reversal succeeded, achieving an accuracy around 70%. This is likely because generating descriptions is more challenging as they have many words and some variety even for the same person. We observe improvement from reverse pre-training in the data-matched case, but not in the compute-matched case. We note that this is perhaps to be expected as the evaluation statements are fictitious and never appeared in the pre-training data.
@@
std. (compute-matched) 65.1 74.4 41.2 47.7 62.7 67.6 32.1 27.0 27.1 49.4
std. (data-matched) 60.5 71.6 41.5 44.5 59.9 64.2 30.0 27.2 27.9 47.5
reverse (token) 63.7 72.9 41.6 45.1 60.0 65.7 30.5 28.0 25.8 48.1
reverse (entity) 62.7 72.3 40.9 45.5 59.4 65.1 29.4 25.4 27.7 47.6
reverse (rand k=25) 63.0 73.2 41.6 46.5 62.0 67.6 31.7 26.4 27.4 48.8
@@ 3.5 Analysis & ablation experiments
In Sections 3.1, 3.2, 3.3 and 3.4 we showed that reverse training helps to mitigate the reversal curse. Here, we explore if our method disrupts zero-shot performance on common evaluation tasks: BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC easy and challenge (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018). We also report 5-shot performance on the aggregated MMLU benchmark (Hendrycks et al., 2020). Evaluation results are summarized in Table 7. We observe that our random reversal model is 1.3 points better than the standard data-matched 1.4B model trained in the standard forward direction, and is only 0.6 points behind the compute-matched model in accuracy on average, despite being trained on half of the tokens. We note that token reversal works slightly better than entity reversal on these standard benchmarks, and both are superior to the data-matched standard training as well.
We also find that reversed models can not only generate text in the normal left-to-right direction, but can generate the beginning of the text given a continuation — a capability that standard models lack (see an example in Appendix B Table 8). This reverse generation function can be useful in itself, for example for instruction backtranslation (Li et al., 2023).
@@
To understand the effect of segment granularity when reversing sequences, we evaluate the performance of the following training methods on the fictitious celebrities task: standard finetuning, token and word reversal finetuning, entity-preserving reversal finetuning, and random segment reversal finetuning with varying $k$ as described in Section 2. The results are summarized in Appendix B Table 9. In general, we find that reversing at a fine-grained level such as token or word level does not significantly help to resolve the reversal curse, and only improves performance on the reverse tasks by 2-3%. Preserving entities during reversal makes it possible to predict names, but not descriptions. This indicates a close relation between the unit of reversal training and the target “concepts” (e.g. names, descriptions) of the reversal task. Similarly, the random segment reversal performs poorly at predicting descriptions when the segment length limit is set lower than the typical length of a description. The results from Section 3.1 also support this hypothesis.
@@ 5 Conclusion
In this paper, we introduced a simple yet effective training method to help remedy the reversal curse in LLMs. Our reverse training works by first segmenting the input sequence into chunks and then reversing the ordering of chunks, but leaves the word-ordering in each chunk intact. A chunk can be a token, a word, an entity name, or a random number of tokens. The model is then trained on both the original sequences, and this reversed data. We evaluated on a symbolic reverse task and a reversing biography task that both demonstrated the necessity of preserving word-ordering within chunks. Next, we applied our reverse training to the realistic setting of LLM pre-training, which minimized the reversal curse on real-world knowledge. Evaluations on common benchmark tasks reveal that reverse training (particularly random segment reversal) during pre-training does not interfere with the forward prediction ability of LLMs, and actually improves metrics in the data-bound (rather than compute-bound) setting compared to standard training. When our method is applied to finetuning on fictitious facts, prediction accuracy rose from 0% to 70-100%. The reversal curse is a serious flaw in how LLMs acquire knowledge and reverse training opens a new promising direction in resolving it.
@@ Appendix B Transformer model pre-training
We train a transformer model with $dim=2048$, $n\_layers=24$, and $n\_heads=16$, resulting in 1.4B parameters. Training data and hyperparameter setup mostly repeats the one from Touvron et al. (2023b). To adapt for the relatively smaller model size, we increase the learning rate to $4.0e{-4}$, and the global batch size was capped at 2M due to the limited number of GPUs. During training, we observe a fixed gap in training perplexity between baseline models and reverse training (Figure 1). The loss of the baseline model is measured on data in the standard direction, while the reverse training loss covers data in both directions. We posit that the reverse training doesn’t interfere with forward learning — thus, the model’s performance does not degrade on standard benchmarks in data-match conditions, and because we observe a match in the convergence rate of the reverse trained models with the baseline model when it’s trained on about 50% of the data.
@@
We also find that reversed models can not only generate text continuations in the normal left-to-right direction, but can generate the beginning of the text given a continuation — a capability that standard models lack. We give an example in Table 8.
Query ”The Star-Spangled Banner” is the $\rightarrow$ $\leftarrow$ national anthem of the United States written by Francis Scott Key.
standard $\rightarrow$ national anthem of the United States. N/A
reverse (token) $\rightarrow$ national anthem of the United States. The Star Spangled Banner is the $\leftarrow$
reverse (entity) $\rightarrow$ national anthem of the United States. The Star-Spangled Banner is the $\leftarrow$
Finally, we evaluate the performance on the fictitious celebrities task in different ablation setups, varying pre-training and finetuning training approaches. Results are summarized in Table 9.
forward reverse forward reverse
Model size: 1.4B
reverse (token) reverse (token) 78.3 0.0 100 2.7
reverse (entity) standard 78.0 0.0 96.3 0.3
reverse (entity) reverse (word) 71.0 2.7 94.7 2.0
reverse (entity) reverse (entity) 77.0 78.3 95.3 2.3
std. (compute-matched) reverse (rand k=5) 77.0 52.3 96.0 10.7
std. (compute-matched) reverse (rand k=10) 74.7 85.3 93.7 33.7
std. (compute-matched) reverse (rand k=25) 77.3 96.3 97.7 70.7
std. (compute-matched) reverse (rand k=50) 77.3 89.3 93.0 67.3